# Shuchang Zhang, Hui Zhang, Hongxia Wang

### v2

@@ 2.2 Implicit Gradient RED
A function $h:\mathbb{R}^{n}\to\mathbb{R}^{n}$ defined everywhere is the proximal operator of a proper convex l.s.c. (lower semicontinuous) function $\phi:\mathbb{R}^{n}\to\mathbb{R}^{+}$ if, and only if the following conditions hold jointly:
@@
    there exists a (convex l.s.c.) function $\psi$ such that for each $\mathbf{x}\in\mathbb{R}^{n},h(\mathbf{x})=\nabla\psi(\mathbf{x})$;
@@
Assume that a denoiser $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ satisfies homogeneity and Assumption (A), and has symmetric Jacobian, then the gradient of RED defines an implicit regularization $\phi:\mathbb{R}^{n}\to\mathbb{R}^{+}$ such that
(11) $R(\mathbf{x})=\nabla g(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x}),D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x}).$
Proof 2.6.
Let $\psi(\mathbf{x})=\frac{1}{2}\langle\mathbf{x},\mathbf{x}-D_{\sigma}(\mathbf{x})\rangle,h(\mathbf{x})=R(\mathbf{x})$ in Proposition 2.4. By Assumption (A), there exist a function denoted $\phi^{*}$($\phi^{*}$ is the conjugate of $\phi$) such that $R(\mathbf{x})=\nabla g(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$. By Moreau decomposition [4, Theorem 6.44], $D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x})$.
@@ 3 Halpern-type preconditioned proximal point (HPPP)
Compared to KM iteration, Halpern iteration [19] is an effective method to find a particular fixed point, i.e.,
(12) $\mathbf{u}^{k+1}=\lambda_{k+1}\mathbf{u}^{0}+(1-\lambda_{k+1})T\mathbf{u}^{k},$
the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ with suitable $\{\lambda_{k}\}_{k\in\mathbb{N}}$ strongly converges to the projection of the initial point $\mathbf{u}^{0}$ to $\mathrm{Fix}(T)$  [19, 26, 40], i.e., $\mathbf{u}^{*}=P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})$, where $P_{\Omega}(\mathbf{u}^{0})=\arg\min_{\mathbf{u}\in\Omega}\left\|\mathbf{u}-\mathbf{u}^{0}\right\|^{2}$ denotes the standard metric projection. Using the PPP to solve the inclusion problem (3), the fixed point of $\mathcal{T}$ is the solution of (3). Choosing an appropriate anchor point, such as the degraded image or an image filled with ones or zeros, can bring the projection closer to the ground-truth solution, resulting in better restoration performance. Hence, the HPPP algorithm (5) theoretically converges to the ground-truth solution under some known anchor point prior $\mathbf{a}$. When $\mathcal{T}$ satisfies the mild condition $\left\|\mathcal{T}\mathbf{u}-\mathcal{T}\mathbf{v}\right\|\leq C\left\|\mathbf{u}-\mathbf{v}\right\|_{\mathcal{M}}(C>0)$ [5], the HPPP algorithm is able to find a unique solution $\mathbf{u}^{*}$ of (3), where $\mathbf{u}^{*}=\arg\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$(see Lemma A.4).
@@ 3.1 Convergence analysis
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz. Let $\{\mathbf{u}^{k}\}$ be the sequence generated by HPPP (5). Assume that every weak cluster point of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$, and $\{\mu_{k}\}_{k\in\mathbb{N}}$ satisfies
@@
Then $\mathbf{u}^{k}$ converges strongly to $\mathbf{u}^{*}$ which is the unique solution of $\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$.
@@
Firstly, we show that $\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\to 0$, where
$\mathbf{u}^{*}=P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})={\arg\min}_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}.$
Then, we have
@@
where $\delta_{k}=\mu_{k}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}^{2}+2\mu_{k}(1-\mu_{k})\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k-1}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$. By Lemma A.2, $\{\mathbf{u}^{k}\}$ is bounded, there exists $\tilde{\mathbf{u}}$ such that $\mathbf{u}^{k}\rightharpoonup\tilde{\mathbf{u}}\in\mathrm{Fix}(\mathcal{T})$ according to the known condition. Assume that $\mathbf{u}^{k_{n}}\rightharpoonup\tilde{\mathbf{u}}$ such that
$\limsup_{k\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathbf{u}^{k}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$ $=\lim_{n\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathbf{u}^{k_{n}}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$
Since $\mathbf{u}^{*}$ is the unique solution of $\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{a}-\mathbf{u}\right\|_{\mathcal{M}}^{2}$ according to Lemma A.4, which solves
$\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\leq 0,\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T}).$
Therefore $\limsup_{k\to\infty}\delta_{k}\leq 0$, from Lemma A.1 we obtain $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}=0$. we then prove $\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|\to 0$.
$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{*}\right\|$ $=\left\|\mu_{k+1}(\mathbf{a}-\mathbf{u}^{*})+(1-\mu_{k+1})(\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\|$
@@
$\leq\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+C(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}},$
it easily follows that $\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|\to 0$ from $\mu_{k+1}\to 0$ and $\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\to 0$.
It is easy to verify that $\mu_{k}=\frac{1}{k^{\alpha}}(0<\alpha\leq 1)$ satisfy conditions (i)-(iii). Under mild conditions(i)-(iii), compared with [2], we have provided an alternative method to obtain the limit point as $\mathcal{M}$-projection of the initial point.
Corollary 3.3.
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be a maximal operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz. Let $\{\mathbf{u}^{k}\}$ be the sequence generated by HPPP (5). If $\{\mu_{k}\}_{k\in\mathbb{N}}$ satisfies $\sum_{k\in\mathbb{N}}\mu_{k}=\infty,\mu_{k}\to 0(k\to\infty),\lim_{k\to\infty}\frac{\mu_{k+1}-\mu_{k}}{\mu_{k}}=0$ or $\sum_{k\in\mathbb{N}}\left|\mu_{k+1}-\mu_{k}\right|<\infty$, then $\mathbf{u}^{k}$ converges strongly to $\mathbf{u}^{*}$, which is the unique solution of $\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$.
@@
Assume that $\mathbf{u}^{k}\rightharpoonup\mathbf{u}$, $\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\to 0$ (Lemma A.2), it follows that
$\mathcal{T}\mathbf{u}^{k}=\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}+\mathbf{u}^{k}\rightharpoonup 0+\mathbf{u}=\mathbf{u},$
i.e., $\mathcal{T}\mathbf{u}^{k}\rightharpoonup\mathbf{u}$. From $\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\to 0$ and $\mathcal{T}\mathbf{u}^{k}=(\mathcal{M}+\mathcal{A})^{-1}\mathcal{M}\mathbf{u}^{k}$ we have
(13) $\mathcal{A}\mathcal{T}\mathbf{u}^{k}\owns\mathcal{M}(\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k})\to 0.$
By the maximality of $\mathcal{A}$ we have that $\mathcal{A}$ is closed in $\mathcal{H}_{weak}\times\mathcal{H}_{strong}$ (see [1, Proposition 20.38]), hence $0\in\mathcal{A}\mathbf{u}$, i.e., $\mathbf{u}$ is a fixed point of $\mathcal{T}$. Thus, every weak cluster point of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$, it holds by Theorem 3.1.
@@
Compared Theorem 3.1, Corollary 3.3 with [5, Theorem 2.9, convergence] and [5, Corollary 2.10], $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by HPPP converges strongly to a particular fixed point of $\mathcal{T}$. All conditions are the same except we only add the mild assumption about $\{\mu_{k}\}_{k\in\mathbb{N}}$ used for convergence analysis. $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ can converges weakly to the same fixed point of $\mathcal{T}$ since $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|\to 0$. The Lipschitz regularity of $(\mathcal{M}+\mathcal{A})^{-1}$ is a mild assumption especially in applications to splitting algorithms, which is used to prove the uniqueness of $\mathcal{M}$-projection and guarantee the boundedness of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$.

### v4

@@ 2.2.1 PnP prior
A function $h:\mathbb{R}^{n}\to\mathbb{R}^{n}$ defined everywhere is the proximal operator of a proper convex l.s.c. (lower semicontinuous) function $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ if, and only if the following conditions hold jointly:
@@
    there exists a convex l.s.c. function $\psi$ such that for each $\mathbf{x}\in\mathbb{R}^{n},h(\mathbf{x})=\nabla\psi(\mathbf{x})$;
@@
2.2.2 RED prior
Romano et al. introduced a famous RED [47], which is defined by
(10) $g_{\text{red}}(\mathbf{x})=\frac{1}{2}\langle\mathbf{x},\mathbf{x}-D_{\sigma}(\mathbf{x})\rangle,$
where $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a denoiser which is assumed to obey the following assumptions,
  1. (C1)
    Local Homogeneity: For all $\mathbf{x}\in\mathbb{R}^{n}$, $D_{\sigma}\left((1+\epsilon)x\right)=(1+\epsilon)D_{\sigma}(\mathbf{x})$ for sufficiently small $\epsilon>0$.
  2. (C2)
    Differentiability: The denoiser $D_{\sigma}(\cdot)$ is differentiable.
  3. (C3)
    Jacobian Symmetry [45]: $[\nabla D_{\sigma}(\mathbf{x})]^{\mathrm{T}}=\nabla D_{\sigma}(\mathbf{x}),\forall\mathbf{x}\in\mathbb{R}^{n}$.
  4. (C4)
    Strong Passivity: The spectral radius of the Jacobian satisfies $\eta\left(\nabla D_{\sigma}(\mathbf{x})\right)\leq 1$.
If $D_{\sigma}(\mathbf{x})$ satisfies C1C2, and C3, then $\nabla g_{\text{red}}(\mathbf{x})=\mathbf{x}-D_{\sigma}(\mathbf{x})=R(\mathbf{x})$. Moreover, if the denoiser satisfies C4, then RED is convex and $D_{\sigma}$ is nonexpansive.
2.2.3 Implicit Gradient RED
The gradient of RED is exactly the denoising residual. We show that under ideal conditions and assuming the residual is nonexpansive, both the denoiser and its residual function as proximal operators of implicit regularizers. Even if all RED conditions are hard to satisfy, learning FNE operators can still achieve this.
The nonexpansive assumption has theoretically played a crucial role in PnP and RED. Ryu et al. proposed the following assumption to guarantee the convergence of PnP methods [49],
(A) $\left\|(I-D_{\sigma})(\mathbf{u})-(I-D_{\sigma})(\mathbf{v})\right\|\leq\varepsilon\left\|\mathbf{u}-\mathbf{v}\right\|,$
for any $\mathbf{u},\mathbf{v}\in\mathbb{R}^{n}$ and $\varepsilon\leq 1$, the real spectral normalization is used to obtain nonexpansive residual. Under the conditons C1C2C3 and C4 of RED and assumption (A), we can further discuss the relationship between $\nabla g_{\text{red}}$ and the proximal operator, and prove that there exists an implicit regularizer $\phi$ such that $D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x})$ and $R(\mathbf{x})=\nabla g_{\text{red}}(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, where $\phi^{*}$ is the conjugate of $\phi$.
@@ 2.2.3 Implicit Gradient RED
Assume that a denoiser $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ satisfies conditons C1C2C3C4 and assumption (A), then there exists an implicit regularizer $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ such that
(11) $D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x}),R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x}).$
Proof 2.7.
By assumption condtions, the RED term $g_{\text{red}}$ is convex and $\nabla g_{\text{red}}=I-D_{\sigma}=R$. Let $\psi(\mathbf{x})=\frac{1}{2}\left\|\mathbf{x}\right\|^{2}-g_{\text{red}}$, then $D_{\sigma}=I-\nabla g_{\text{red}}=\nabla\psi$, which is a GS denoiser. Since $R$ is nonexpansive and $D_{\sigma}$ is differentiable, then $\eta(\nabla^{2}\psi)=\eta(\nabla D_{\sigma})=\eta(I-\nabla R)\in[0,1]$, it follows that $\psi$ is convex. Applying Proposition 2.5 together with $h(\mathbf{x})=D_{\sigma}(\mathbf{x}),\psi=\frac{1}{2}\left\|\mathbf{x}\right\|^{2}-g_{\text{red}}$ and Moreau decomposition [5, Theorem 6.44], i.e., $\mathrm{prox}_{\phi}(\mathbf{x})+\mathrm{prox}_{\phi^{*}}(\mathbf{x})=I$, then there exist a function $\phi$ and its conjugate $\phi^{*}$ such that $D_{\sigma}(\mathbf{x})=\nabla\psi(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x})$, and $R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, which completes the proof.
Satisfying all RED conditions is challenging; in practice, only the nonexpansiveness assumption is required. From the perspective of monotone operator theory, Lemma 2.8 [2, Corollary 23.9] provides a necessary and sufficient condition for a denoiser to be a proximal operator of an implicit regularizer.
Lemma 2.8.
The mapping $T:\mathcal{X}\to\mathcal{X}$ is the resolvent of a maximal monotone operator if and only if $T$ is FNE.
Moreover, by [2, Proposition 4.4], the operator $T$ is FNE if and only if there exists a nonexpansive operator $S:\mathcal{X}\to\mathcal{X}$ such that $T=\frac{S+I}{2}$. Therefore, the core issue in explaining denoiser prior as a proximal operator of a convex regularizer is training a nonexpansive operator. Recent developments, such as spectral normalization [49, 21], learning maximally monotone operators [42], learning FNE denoisers [54, 10], and learning pseudo-contractive denoisers [57], have made the nonexpansive assumption more realistic in practice. Constructing the FNE denoiser $D_{\sigma}$ is feasible, and similarly, the residual $R=I-D_{\sigma}$ can also be FNE in (7) with $\mathcal{M}=I$. They can both be the proximal operators of implicit regularizers.
@@ 3 Halpern-type preconditioned proximal point (HPPP)
Compared with KM iteration, the classic Halpern iteration offers the advantage of strong convergence in infinite-dimensional Hilbert spaces [29]. The Halpern iteration [26]
$\mathbf{u}^{k+1}=\lambda_{k+1}\mathbf{u}^{0}+(1-\lambda_{k+1})T\mathbf{u}^{k}$
is an implicitly regularized method to find a particular fixed point [22], and the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by the Halpern iteration with suitable $\{\lambda_{k}\}_{k\in\mathbb{N}}$ strongly converges to the projection $P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})$ [26, 37, 58, 30, 61, 44, 29], where $P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})=\arg\min_{\mathrm{Fix}(T)}\left\|\mathbf{u}-\mathbf{u}^{0}\right\|^{2}$. When PPP is used to solve the inclusion problem (3), the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by PPP can only converge weakly to some uncertain fixed point of $\mathcal{T}$. Based on the Halpern iteration, we propose the Halpern-type preconditioned proximal point (HPPP) algorithms (5).
@@ 3.1 Convergence analysis
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz. Let $\{\mathbf{u}^{k}\}$ be the sequence generated by HPPP (5). Assume that every weak cluster point of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$, and $\{\mu_{k}\}_{k\in\mathbb{N}}$ satisfies
@@
Then $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ converges strongly to $\mathbf{u}^{*}$.
@@
First, we show $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}=0$. Based on the definition of the upper limit, we take a subsequence $\{\mathbf{u}^{k_{n}}\}_{n\in\mathbb{N}}$ of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ such that
(12) $\limsup_{k\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}=\lim_{n\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}.$
By Lemma A.2(i), the sequence $\{\mathbf{u}^{k_{n}}\}_{n\in\mathbb{N}}$ is bounded, there exists a weakly convergent subsequence $\{\mathbf{u}^{k_{n_{j}}}\}_{j\in\mathbb{N}}$, we may assume $\mathbf{u}^{k_{n_{j}}}\rightharpoonup\tilde{\mathbf{u}}$. According to the conditions of the theorem, we have $\tilde{\mathbf{u}}\in\mathrm{Fix}(\mathcal{T})$.
Applying Lemma A.2 together with conditions (i)-(iii), we obtain $\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\to\mathbf{0}$, then $\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\rangle_{\mathcal{M}}\to 0$. Hence, we have
$\lim_{n\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$ $=\lim_{j\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},(\mathcal{T}\mathbf{u}^{k_{n_{j}}}-\mathbf{u}^{k_{n_{j}}})+(\mathbf{u}^{k_{n_{j}}}-\mathbf{u}^{*})\right\rangle_{\mathcal{M}}$
$=\lim_{j\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathbf{u}^{k_{n_{j}}}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$
By Lemma B.1, we have
(13) $\limsup_{k\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}=\left\langle\mathbf{a}-\mathbf{u}^{*},\tilde{\mathbf{u}}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\leq 0.$
Denote $\delta_{k+1}=\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}^{2}+2(1-\mu_{k+1})\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$, notice that
@@
applying Lemma A.1 together with conditions (i)-(ii) and $\limsup_{k\to\infty}\delta_{k}\leq 0$, then we have $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}=0.$
By the definition of $\mathbf{u}^{k+1}$ we immediately obtain
$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{*}\right\|$ $=\left\|\mu_{k+1}(\mathbf{a}-\mathbf{u}^{*})+(1-\mu_{k+1})(\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*})\right\|$
@@
(14) $\leq\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+C(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}},$
where the last inequality follows from (26) and $C$ is defined in (25). Since $\lim_{k\to\infty}\mu_{k}=0$ and $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}=0$, and thus combined with (14) yields that
$\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|=0,$
which ends the proof of Theorem 3.1.
It is easy to verify that $\mu_{k}=\frac{1}{k^{\alpha}}(0<\alpha\leq 1)$ satisfes conditions (i)-(iii) of Theorem 3.1. Compared with [3], we have provided an alternative method to obtain the limit point as $\mathcal{M}$-projection of the initial point. The work [3] requires that $\mathcal{A}$ be linear, whereas we consider a more general case for $\mathcal{A}$. By introducing the maximal monotonicity assumption on $\mathcal{A}$, we can prove that every weak cluster point $\mathbf{u}$ of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$. According to Theorem 3.1, we obtain the following Proposition 3.3.
Proposition 3.3.
If the conditions of Theorem 3.1 hold, and $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ is a maximal monotone operator, then $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ converges strongly to $\mathbf{u}^{*}$.
@@
Assume that $\mathbf{u}^{k_{n}}\rightharpoonup\mathbf{u}$. By Lemma A.2(iii), we have $\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{k_{n}}\to\mathbf{0}$, it follows that
$\mathcal{T}\mathbf{u}^{k_{n}}=(\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{k_{n}})+\mathbf{u}^{k_{n}}\rightharpoonup\mathbf{0}+\mathbf{u}=\mathbf{u},$
i.e., $\mathcal{T}\mathbf{u}^{k_{n}}\rightharpoonup\mathbf{u}$. Using $\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{k_{n}}\to\mathbf{0}$ we have
$\mathcal{A}\mathcal{T}\mathbf{u}^{k_{n}}\owns\mathcal{M}(\mathbf{u}^{k_{n}}-\mathcal{T}\mathbf{u}^{k_{n}})\to\mathbf{0}.$
Since $\mathcal{A}$ is maximal, we have that $\mathcal{A}$ is closed in $\mathcal{H}_{weak}\times\mathcal{H}_{strong}$ [2, Proposition 20.38], hence $\mathbf{0}\in\mathcal{A}\mathbf{u}$, i.e., $\mathbf{u}\in\mathrm{zer}\mathcal{A}=\mathrm{Fix}(\mathcal{T})$. Thus, we prove that every weak cluster point of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$. This completes the proof of Proposition 3.3.
@@
Compared Theorem 3.1, Proposition 3.3 with [9, Theorem 2.9] and [9, Corollary 2.10], where the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by HPPP converges strongly to a particular fixed point of $\mathcal{T}$. All conditions are the same except for additional assumption about $\{\mu_{k}\}_{k\in\mathbb{N}}$. The Lipschitz regularity of $(\mathcal{M}+\mathcal{A})^{-1}$ is a mild assumption especially in applications to splitting algorithms, which is used to prove the uniqueness of $\mathcal{M}$-projection and guarantee the boundedness of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$.
Hundal constructed an example in which the alternating projection $(P_{U}P_{V})^{k}\mathbf{x}^{0}$ converges weakly but not strongly, where $U$ is a closed convex cone and $V$ is a closed hyperplane [31]. Building on Hundal’s counterexample, [13] showed that the weak convergence of the DRS algorithm cannot be improved to strong convergence. Similarly, using the same counterexample, we can demonstrate that PPP also fails to converge strongly.
Counterexample 1.
Suppose that $\mathcal{H}$ is infinite-dimensional and separable. Let $V$ and $U$ be Hundal’s hyperplane and Hundal’s cone, respectively, and set
$\mathcal{A}_{1}:\mathbf{x}\mapsto\begin{cases}V^{\perp}&\text{ if }\mathbf{x}\in V,\\ \emptyset&\text{ if }\mathbf{x}\in V,\end{cases}\text{ and }\mathcal{A}_{2}=(P_{V}\circ P_{U}\circ P_{V})^{-1}-I.$
Then the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by PPP (8) with $\lambda_{k}=1$ converges weakly, but not strongly, to a zero of $\mathcal{A}_{1}+\mathcal{A}_{2}$.
Proof 3.6.
As proved in [13], the operators $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are maximally monotone, and $\mathbf{0}\in\mathrm{zer}(\mathcal{A}_{1}+\mathcal{A}_{2})$, and the following sequence $\{\mathbf{w}^{k}\}_{k\in\mathbb{N}}$
(15) $\mathbf{w}^{k+1}=\mathbf{w}^{k}+J_{\mathcal{A}_{2}}(2J_{\mathcal{A}_{1}}\mathbf{w}^{k}-\mathbf{w}^{k})-J_{\mathcal{A}_{1}}\mathbf{w}^{k}$
generated by the DRS algortihm can weakly converges to $\mathbf{0}$ and $\mathbf{w}^{k}\not\rightarrow\mathbf{0}$.
Since the above (15) can be viewed as a special PPP with
$\mathcal{A}=\begin{pmatrix}\mathcal{A}_{1}&I\\ -I&\mathcal{A}_{2}\end{pmatrix},\mathcal{M}=\begin{pmatrix}I&-I\\ -I&I\end{pmatrix},$
and $\mathbf{w}^{k}=\mathcal{C}^{*}\mathbf{u}^{k}=\mathbf{x}^{k}-\mathbf{y}^{k}$, where $\mathcal{C}^{*}=(I,-I)$. By [9, Corollary 2.15], $\mathbf{u}^{k}\rightharpoonup\mathbf{0}$. Indeed, if by contraction $\mathbf{u}^{k}\to\mathbf{0}$, then
$\left\|\mathbf{w}^{k}-\mathbf{0}\right\|=\left\|\mathcal{C}^{*}\mathbf{u}^{k}-\mathbf{0}\right\|\leq\left\|\mathcal{C}\right\|\cdot\left\|\mathbf{u}^{k}-\mathbf{0}\right\|\to 0,$
which contradicts the fact $\mathbf{w}^{k}\not\rightarrow\mathbf{0}$. Thus $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ can only converge weakly to $\mathbf{0}$.
Remark 3.7.
From Counterexample 1, HPPP has a strong convergence advantage over PPP in infinite-dimensional spaces, which has important theoretical value.
Remark 3.8.
Regarding the choice of the anchor point, we can use the degraded observed image, the denoised image, or the restarted technique (see Algorithm 2). The restarted technique is an effective method in machine learning [41] and LP [17, 38].
Algorithm 2 Restarted HPPP
1:Initialization $\mathbf{u}^{0,0}=\mathbf{a}^{0}$, total iteration $N>0$, period $q>0$, and epoch $N_{e}=\lfloor\frac{N}{q}\rfloor$
2:for $n=1,2,\cdots,N_{e}$ do
3:     for $k=0,1,\cdots,q-1$ do
4:         $\mathbf{u}^{n,k+1}=\mathrm{HPPP}(\mathbf{u}^{n,k},\mathbf{a}^{n})$
5:     end for
6:     $\mathbf{a}^{n+1}=\mathbf{u}^{n,q}$
7:end for
8:$\mathbf{u}^{N_{e},q}$